OpenStudy (anonymous):

John pushes his brother Danny on a skateboard. John applies a force of 39 N and Danny's acceleration is 0.9 m/s2. What is the combined mass of Danny and his skateboard? A. 43 kg B. 35 kg C. 31 kg D. 48 kg

OpenStudy (anonymous):

\[Force=mass*acceleration\]

OpenStudy (anonymous):

39/0.9=43

OpenStudy (shamim):

U know F=ma 39=m*0.9 m=?

rvc (rvc):

@tamela12 u just need to to the formula then put in the values and get the answer :)